OpenStudy (anonymous):

solve the equation \[\sqrt{3w+4}+w=8\] im not sure what to do with the extra w in this.

OpenStudy (anonymous):

\[\sqrt{3w+4}=w+8\]

OpenStudy (anonymous):

\[(\sqrt{3w+4)^2}+w=8^2\]

OpenStudy (anonymous):

i might be going the wrong way help lol

OpenStudy (lostinmath101):

Subtract the w to the other side.

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{3w+4}+w=8\] \[\Large \sqrt{3w+4}=8-w\] \[\Large \left(\sqrt{3w+4}\right)^2=(8-w)^2\] \[\Large 3w+4=(8-w)^2\] Do you see where to go from here?

OpenStudy (lostinmath101):

So now it's\[\sqrt{3w+4}=w+8\]

OpenStudy (lostinmath101):

square both sides So: 3w=4=(w+8)^2

OpenStudy (radar):

Did you keep track of your signs?

OpenStudy (anonymous):

im trying to

OpenStudy (lostinmath101):

Now we have 3w+4=\[w ^{2}+16w +64]

OpenStudy (lostinmath101):

Subtract 3w from both sides: 4= w^2+13w+64 Then subtract 4 from both sides: 0= \[w ^{2}+13w+68\]

OpenStudy (radar):

\[(8+w)^{2}\neq(8-w)^{2}\]

OpenStudy (anonymous):

shouldnt be a - should there? the original is all +

OpenStudy (lostinmath101):

They're both postives, so the numbers in the parentheses can be positve or negative. However 13w is positive, so it's positive in both. Mistake: it's + 60

OpenStudy (anonymous):

64-4=60

OpenStudy (lostinmath101):

\[w ^{2}+13w+60=0\]

OpenStudy (anonymous):

what you do to one side you do to the other rule got it, i got that answer with you, i was just lost on the extra w and how to input it into everything

OpenStudy (lostinmath101):

I'm sorry i followed the wrong ex. It's (8-w)^2

OpenStudy (anonymous):

the original was \[\sqrt{3w+4}+w=8\]

OpenStudy (lostinmath101):

Okay, so \[-w ^{2}-16w+64\]

jimthompson5910 (jim_thompson5910):

no, (8-w)^2 = w^2 - 16w + 64

OpenStudy (lostinmath101):

then subtract 3w and 4 to get \[-w ^{2}-19w+60\]. Carry it over to the other side so it equals \[w ^{2}+19w-60\]

OpenStudy (anonymous):

<~~~~is really confused now lol

OpenStudy (lostinmath101):

Me too..

jimthompson5910 (jim_thompson5910):

It's probably best to start over then

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

\[\sqrt{3w+4}+w=8\] original equation

OpenStudy (radar):

\[\sqrt{3w+4}=8-w\]\[3w+4=(8-w)^{2}\]\[3w+4=64-16w+w ^{2}\]\[w ^{2}-19w+60=0\]

OpenStudy (radar):

Use quadratic formula to solve.

OpenStudy (radar):

Or factor.

OpenStudy (radar):

(w-15)(w-4)=0

OpenStudy (radar):

w=15 w=4 Check in original to verify.

OpenStudy (anonymous):

ok im following i had to re-write so i wasnt getting lost

OpenStudy (radar):

Good, factoring is always easier than using the quadratic, provided it is factorable lol.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

omg ty, i so got lost there for a few